YOUR REMOTE CONTROL HAS BUTTONS FOR DIGITS 0-9. IF YOU PRESS TWO BUTTONS, HOW MANY NUMBERS ARE POSSIBLE IF
A) THE SAME BUTTONS MAY BE PRESSED TWICE
B) THE SAME BUTTONS MAY NOT BE PRESSED TWICE

Answer by Alan3354(69443)

YOUR REMOTE CONTROL HAS BUTTONS FOR DIGITS 0-9. IF YOU PRESS TWO BUTTONS, HOW MANY NUMBERS ARE POSSIBLE IF
A) THE SAME BUTTONS MAY BE PRESSED TWICE
1 of 10, then 1 of 10 = 10*10 = 100
--------------------
B) THE SAME BUTTONS MAY NOT BE PRESSED TWICE
1 of 10, then 1 of 9 = 10*9 = 90
------------
PS Don't use all CAPS, it irritates people.